An equimolar solution of two volatile liquids A of M= 150 g/mol and B(M=200 g/mol)
has vapour pressure 500 torr.If vapour pressure of pure B is 400 torr the fraction of component
A in vapour phase is

Grade:12

1 Answers

Vikas TU
7 years ago
Total pressure = Xa*Pa + Xb*Pb
500 = 0.5*Pa  + 0.5*400
Pa/2 = 500 – 200 = 300
Pa = 600 torr. is the vapor pressure for pure B.